# Vision-Language Models for Medical Report Generation and Visual Question Answering: A Review

### v1

@@ 7 Challenges & Future Directions
In medical AI, the future holds great promise and, concurrently, poses notable challenges [Aco+22]. As technology advances, the integration of VLMs into the healthcare sector has the potential to revolutionize diagnostics, treatment planning, and patient care. Future medical VLMs may offer enhanced capabilities in understanding complex clinical scenarios, generating detailed reports of medical images, and facilitating seamless communication between healthcare professionals and AI systems. However, these advancements come with challenges.
@@
A significant challenge in developing effective medical VLMs is the lack of diverse and representative medical datasets. This limitation restricts the comprehensive training of VLMs, impeding their ability to understand the complexities of diverse and rare clinical scenarios [Moo+23]. Retrieval-augmented generation (RAG) presents a potential solution by enhancing training datasets through incorporation of retrieved relevant information. While RAG usually involves a frozen model during training, exploring the pre-training of VLMs within the RAG framework opens up a new avenue of research [Zha+23c]. This innovative approach could potentially enhance the robustness of VLMs, especially in handling new and unforeseen medical cases.
@@
Traditional metrics may fall short in capturing the nuanced complexities of clinical language, posing a barrier to reliable evaluations of VLM performance [Yu+23]. This issue becomes particularly evident when evaluating the accuracy of medical reports or addressing open-ended medical queries, where metrics need to discern clinically relevant distinctions. Therefore, development and adoption of specialized metrics tailored for medical RG and VQA become imperative. Such metrics are pivotal not only for evaluating model performance but also for assessing aspects like generalization, efficiency, and robustness. The establishment of these metrics will significantly contribute to fostering precise evaluations and continual advancements in the capabilities of medical VLMs.
@@
The issue of hallucinations in VLMs poses a significant challenge to their reliability and practical application [Liu+24]. Hallucinations refer to instances where VLMs generate outputs that are not grounded in the provided images, or are inconsistent with the established knowledge. In medical contexts, these hallucinations can have serious consequences, leading to inaccurate diagnostic information or treatment recommendations. One identified root cause of hallucinations is the lack of alignment between visual and textual modalities [Sun+23]. Training VLMs to effectively align these modalities is crucial in mitigating the risk of hallucinations. For instance, LLaVA-RLHF [Sun+23] achieved hallucination reduction by incorporating RLHF to align different modalities.
@@
Overcoming catastrophic forgetting poses an additional challenge in the development of medical VLMs. Catastrophic forgetting occurs when a model learning new information inadvertently erases or distorts previously acquired knowledge, potentially compromising its overall competence. Striking a balance during fine-tuning can be crucial; moderate fine-tuning can be helpful to adapt the model to a specific task, while excessive fine-tuning can lead to catastrophic forgetting [Zha+23]. Leveraging methodologies from continual learning [Wan+23, Zho+23a, CR24] might be useful in the context of medical VLMs, where the ability to adapt and accumulate knowledge across diverse clinical tasks is paramount. Continual learning focuses on training models to sequentially learn from and adapt to new data over time while retaining knowledge from previously encountered tasks. Also, incorporating adapters within the framework of continual learning can be a valuable tool in mitigating catastrophic forgetting [Zha+23b].
@@
Finally, clinical validation and adoption of VLMs necessitate a collaborative bridge between medical experts and machine learning researchers. Trust, alignment with clinical needs, and ethical deployment are critical components for the successful integration of these models into healthcare workflows. Establishing robust collaborations ensures a dynamic synergy, combining domain expertise with technological advancements. This synergy is essential for the responsible and effective deployment of medical VLMs in healthcare.

### v2

@@ 5 Challenges and Potential Future Directions
In medical AI, the future holds great promise and, concurrently, poses notable challenges [acosta2022biomedicalai]. As technology advances, the integration of VLMs into the healthcare sector has the potential to revolutionize diagnostics, treatment planning, and patient care. Future medical VLMs may offer enhanced capabilities in understanding complex clinical scenarios, generating detailed reports of medical images, and facilitating seamless communication between healthcare professionals and AI systems. However, these advancements come with challenges.
@@
A significant challenge in developing effective medical VLMs is the limited availability of ML-ready diverse and representative medical datasets. This limitation restricts the comprehensive training of VLMs, impeding their ability to understand the complexities of diverse and rare clinical scenarios [moor2023medflamingo]. VLMs with large context windows and RAG present a potential solution by increasing the model’s context through the incorporation of retrieved relevant information. While RAG usually involves a frozen model during training, exploring the pre-training of VLMs within the RAG framework opens up a new avenue of research [zhao2023multimodalrag]. This innovative approach could potentially enhance the robustness of VLMs, especially in handling new and unforeseen medical cases. Furthermore, the pressing concerns surrounding patient data privacy highlight the need for innovative solutions, like federated learning (FL). FL offers a promising strategy to alleviate the scarcity of medical data while prioritizing patient privacy [zhang2021fedlearning]. In this decentralized learning method, models are trained across multiple institutions, and only model weights are shared, not the data. Thus, it effectively addresses major concerns about patient privacy while enabling collaborative model training across diverse datasets.
@@
Traditional metrics may fall short in capturing the nuanced complexities of clinical language, posing a barrier to reliable evaluations of VLM performance [yu2023radgraph]. This issue becomes particularly evident when evaluating the accuracy of medical reports or addressing open-ended medical queries, where metrics need to discern clinically relevant distinctions. Therefore, the development and adoption of specialized metrics tailored for medical RG and VQA is imperative. Such metrics are pivotal not only for evaluating model performance but also for assessing aspects like generalization, efficiency, and robustness. Establishing these metrics will significantly contribute to fostering precise evaluations and continual advancements in the capabilities of medical VLMs.
@@
The issue of hallucinations in generative VLMs poses a significant challenge to their reliability and practical application [liu2024vlmhallucination]. Hallucinations refer to instances where VLMs generate outputs that are not grounded in the provided images or are inconsistent with the established knowledge. In medical contexts, these hallucinations can have serious consequences, leading to inaccurate diagnostic information or treatment recommendations. One identified root cause of hallucinations is the lack of alignment between visual and textual information [sun2023llavarlhf]. Training VLMs to effectively align these data modalities is crucial in mitigating the risk of hallucinations. For instance, LLaVA-RLHF [sun2023llavarlhf] achieved hallucination reduction by incorporating RLHF to align different modalities. Further research is needed into building medical VLMs that base their generation on the factual medical knowledge of minimal hallucination.
@@
Overcoming catastrophic forgetting poses an additional challenge in the development of medical VLMs. Catastrophic forgetting occurs when a model learning new information inadvertently erases or distorts previously acquired knowledge, potentially compromising its overall competence. Striking a balance during fine-tuning can be crucial; moderate fine-tuning can be helpful to adapt the model to a specific task, while excessive fine-tuning can lead to catastrophic forgetting [zhai2023catasforgettingfinetun, khan2023importance]. Leveraging methodologies from continual learning [wang2023contlearning, zhou2023proof, cai2024dynamictransformer, khan2023importance, khan2024brain] might be useful in the context of medical VLMs, where the ability to adapt and accumulate knowledge across diverse clinical tasks is paramount. Continual learning focuses on training models to sequentially learn from and adapt to new data over time while retaining knowledge from previously encountered tasks [khan2024brain]. Also, incorporating adapters within the framework of continual learning can be a valuable tool in mitigating catastrophic forgetting [zhang2023adapterincontlearning].
@@
Finally, clinical validation and adoption of VLMs necessitate a collaborative bridge between medical experts and AI/ML researchers. Trust, alignment with clinical needs, and ethical deployment are critical components for successfully integrating these models into healthcare workflows. Establishing robust collaborations ensures a dynamic synergy, combining domain expertise with technological advancements. This synergy is essential for the responsible and effective deployment of medical VLMs in healthcare.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was partly supported by NSF awards 1903466, 2234836, and 2234468.
[…]
[title=References]